OpenStudy (anonymous):

Find a four digit number so that if a decimal point is placed between the hundreds digit, the resulting number is the average of the two-digit whole number on either side of the decimal point.

OpenStudy (anonymous):

Consider this:\[ d_1d_2d_3d_4 \]

OpenStudy (anonymous):

IT is equal to: \[ d_110^4+d_210^2+d_310^1+d_410^0 \]

OpenStudy (anonymous):

\[ avg=\frac{(10d_1+d_2) +(10d_3+d_4)} 2 \]

OpenStudy (anonymous):

\[ avg=10d_1+d_2+\frac 1{10}d_3+\frac 1{100}d_4 \]

OpenStudy (anonymous):

That's a pretty damn good start.

OpenStudy (anonymous):

You could make it easier and just deal with two digits at a time rather than one digit at a time.

OpenStudy (anonymous):

\[ n=10d_1+d_2\\ m=10d_3+d_4 \]

OpenStudy (anonymous):

\[ \frac{n+m}2=n+\frac m{100} \]

OpenStudy (anonymous):

It will give you multiple solutions, but then you must bound them by \(10\leq n,m\leq 99\)